# Deep Learning Meets Mechanism Design: A Survey

### v2

@@ I Introduction
How do you realize social goals in a modern society consisting of self-interested individuals or agents? This is a question facing all social planners and organisations all the time. If we solve this problem satisfactorily, it has far-reaching implications for creating robust sociological institutions, as well as, for solving numerous modern problems arising at the interface of computer science and microeconomics. Mechanism Design is an important sub-field of economic theory that strives to answer the above question. Mechanism design can be thought of as the reverse engineering of a game to achieve specific desired objectives. In mechanism design, based on the objectives to be achieved, the mechanism induces a game in a way that the desired, specific objectives are realized in an equilibrium of the game [BORGERS15, NISAN07, NARAHARI14, NARAHARI09]. Auctions and matching market protocols are important and popular examples of mechanisms. In this paper, we focus on auctions; however, the results presented apply to more general mechanisms as well.
@@ I-A Motivation for the Survey
Satisfying all of the properties listed above simultaneously is clearly a tall order. Mechanism design theory [NARAHARI14, MASCOLELL95, KRISHNA09] is replete with celebrated results which assert that these properties cannot be simultaneously satisfied. This has led to several methodologies to surmount this limitation. One of the recently emerged, novel methodologies is an approach based on deep learning (DL). The goal of the current survey paper is to discuss this recent approach, based on Deep Learning, for designing mechanisms that achieve maximal subsets of these objectives, approximately, with as little compromise (or regret) as possible. The DL-based approach achieves this by learning a mechanism that approximately satisfies the required set of properties by minimizing a suitably defined loss function. Additionally, the DL-based approach may reduce the computational complexity of deducing the mechanism. In this paper, we present key technical details of the DL-based approach for mechanism design and provide an overview of the key methods that have emerged in this topic.
@@ I-B Outline of the Paper
The paper by Zhang [ZHANG21] is the lone existing survey paper in this area. The paper by Zhang only provides an overview of how deep learning techniques are being applied to the field of online auction mechanism design. The survey explores only a few deep learning methods for designing auctions and serves only a limited purpose. Covers only papers until 2020. Our survey is comprehensive: (a) provides an overview of mechanism design with a detailed illustrative example; (b) covers four broad categories of these mechanisms: revenue maximizing (cost minimizing) mechanisms; welfare maximizing mechanisms; mechanisms with fairness in allocation; and mechanisms with budget balance; and (c) discusses three illustrative applications.
@@ II Mechanism Design: An Overview
Mechanism design can be described as the art of designing the rules of a game to achieve a specific desired outcome [NARAHARI14, MASCOLELL95]. The main focus of mechanism design is to create institutions or protocols that satisfy certain desired objectives, assuming that individual agents are rational and intelligent and will act strategically and may hold private information that is relevant to the decision at hand. In a mechanism design setting, a policy maker (also called social planner or mechanism designer) faces the problem of aggregating the announced preferences of multiple agents into a collective (or social), system-wide decision when the actual preferences of the agents are not publicly known. The essential technique that mechanism design uses to accomplish this is to induce a game among the strategic agents in such a way that in an equilibrium of the induced game, the desired system-wide solution is implemented. Informally, a mechanism induces strategic players to do what the social planner would like them to do. Mechanism design has widespread and significant applications in modern web-based and AI systems [BORGERS15, NISAN07, NARAHARI14, NARAHARI09].
@@ II-A Mechanisms
The notion of mechanisms was first introduced by Leonid Hurwicz (Nobel Laureate in Economic Sciences in 2007) in 1960 [NARAHARI14]. William Vickrey (Nobel Laureate in Economic Sciences in 1996) authored a seminal paper in 1961 [VICKREY61] where he introduced the celebrated Vickrey auction (second price auction) and showed that the mechanism induces buyers in an auction to bid their true valuations without having to worry about what the other buyers bid. The Vickrey auction can be considered as a significant landmark in mechanism design theory. Edward Clarke [CLARKE71] and Theodore Groves [GROVES73] came up with generalizations of Vickrey mechanisms and paved the way for the famous VCG mechanisms (Vickrey-Clarke-Groves) that define a broad class of mechanisms satisfying certain very desirable properties. The discipline of mechanism design has made phenomenal advances in the decades following the 1960s and has found widespread applicability in a wide variety of applications and areas. The scope of mechanisms includes auctions and markets, trading institutions, regulation and auditing, social choice theory, resource allocation in multi-agent systems, electronic commerce, Internet advertisement, crowdsourcing, industrial procurement, spectrum allocation, etc. [BORGERS15, NISAN07, NARAHARI14, NARAHARI09].
Mechanisms in which monetary transfers are allowed (for example, auctions) are known as mechanisms with money whereas mechanisms where there are no monetary transfers (such as voting mechanisms) are called mechanisms without money [NARAHARI14]. In this paper, we are concerned with mechanisms with money. In particular, we focus on auctions, which are arguably the most popular example of mechanisms with money. Although our discussion is centered on auctions, all the results presented here can be generalized to all mechanisms with money.
@@ II-B Auctions
Auctions provide a natural example of mechanisms with money. An auction is a mechanism for assigning a set of goods to a set of agents based on bids announced by the agents. Auctions have been used for hundreds of years for trading of goods and services and their adoption in electronic commerce has risen exponentially in the past few decades. There is intense interest in the design of auctions due to their widespread applicability to traditional settings such as paintings and memorabilia as well as their popular adoption in more recent areas such as Internet advertising, keyword auctions on search engines, airport slot allocation, spectrum allocation, industrial procurement, and resource allocation in general. A survey of auctions appears in several classic articles, including [KLEMPERER99, MCAFEE87].
@@ II-C Desirable Properties of a Mechanism
We now provide a list of important properties that a mechanism is expected to satisfy. This list is taken from [NARAHARI14]. As already stated, a mechanism induces a game among strategic agents in such a way that desirable properties are exhibited in an equilibrium of the game. The different equilibrium solutions that could be considered are dominant strategy equilibrium, Bayesian Nash equilibrium, and ex-post Nash equilibrium. For ease of understanding, we use the example of auctions while describing desirable properties.
@@
Individual rationality ensures that all the agents obtain non-negative utility by participating in the auction. The most powerful version of IR is ex-post IR, which implies that the utility to each participating player will be non-negative irrespective of the bids of the other players. Other forms of IR are interim IR and ex-ante IR [NARAHARI14, MASCOLELL95].
@@ II-D Possibilities and Impossibilities in Mechanism Design
The trade-offs involved in designing mechanisms with desirable properties such as incentive compatibility, individual rationality, social welfare maximization, and budget balance are quite complex and intricate [NARAHARI14, MASCOLELL95]. There is vast literature in mechanism design theory that comprises possibility results and impossibility theorems. As an example, VCG mechanisms satisfy dominant strategy incentive compatability, social welfare maximization, weak budget balance, and individual rationality. They do not satisfy strict budget balance. In fact, it is impossible for any mechanism to satisfy dominant strategy incentive compatability, social welfare maximization, and strong budget balance. Similarly, no mechanism can satisfy incentive compatability, social welfare maximization, strong budget balance, and individual rationality. When we include fairness of allocation and revenue maximization (or cost minimization), the situation gets worse and most subsets of these properties become infeasible.
@@ III An Illustrative Example: Learning a Revenue Maximizing Auction
In this section, we provide the example of an auction mechanism which is a generalization of the famous Myerson auction [MYERSON81] for selling a single indivisible item. Such auctions are relevant in a wide variety of resource allocation settings. While the Myserson auction maximizes the expected revenue subject to Bayesian incentive compatibility and interim individual rationality, the extension we discuss does not satisfy these properties simultaneously and we show how the deep learning based approach is used to tackle this problem.
@@ III-B Ad-Auction: Problem Formulation
Further, designing an auction mechanism that satisfies all the desirable properties simultaneously is a challenging task and it fails in certain scenarios that are described by impossibility results [GIBBARD73],[ARROW12].
@@
The revelation principle [MCAFEE87] considers one shot revelation of true valuations of buyers to the seller by designing an IC mechanism. But it is not the case when players are selfish and may not be truthful for their personal benefits. In this scenario, a seller may be able to extract truthful valuations via multiple communications with the buyers. As opposed to the design of optimal auctions with unlimited communication [MYERSON81], the authors of [MOOKHERJEE14] propose a feasible auction design that maximizes the utility subject to communication feasibility constraints. Here communication feasibility specifies either delay or cost of sending the bid in terms of number of bits. As the number of items $m$ and the number of buyers $n$ increases, the cost of communicating bids or delay in communicating bids also increases. It may also take more iterative communication cycles to extract truthful information about the valuations from the buyers. The authors obtained strong results for designing optimal decentralized decision making under the absence of trade-offs between incentives and information efficiency. In [ZANDT07], the authors study the relation between IC constraints and communication constraints. They also show that there exists a separation between checking IC constraints for Social Choice Function and the design of communication efficient protocol to compute the Social Choice Function for ex post Nash Implementation.
We can use the RegretNet architecture [DUTTING23] to solve the $m$ items and $n$ unit demand buyers problem as shown in Fig. 1. This architecture aims to maximize the expected revenue to the seller under IC. Further details about the architecture are provided in Section IV.
Figure 1: RegretNet: The allocation network for $n$ unit demand buyers with $m$ items (Figure reproduced from [DUTTING23])
@@ IV Learning Revenue Optimal Mechanisms
In [DUTTING23], the authors propose an optimal auction design via RochetNet and RegretNet architectures that satisfy Dominant Strategy Incentive Compatibility (DSIC) and IR. The RochetNet architecture is based on constraining the search space so as to achieve DSIC and in this method the DSIC characterization is encoded within the Neural Network (NN) architecture. This architecture is suitable for a single buyer seeking to purchase multiple items from the auctioneer. In the case of the RegretNet architecture, approximate DSIC is achieved by differential approximation and Lagrangian augmentation. This method entails searching throughout a larger parameter space and is suitable for multi-buyer and multiple-item settings. The authors of [FENG18] propose a DL based revenue optimal auction with private budget constraints of the buyers. This work is an extension of the RegretNet framework proposed in [DUTTING23] that can handle both Bayesian IC and Conditional IC constraints. The authors in [DENG20] introduce a new metric that quantifies Incentive Compatibility in auctions for both static and dynamic environments. This metric is data driven and is evaluated on the ad auction data.
@@
The authors in [CURRY22] design revenue maximizing auctions while approximating incentive compatibility. They build on the RegretNet architecture by introducing a differentiable matching approach in the allocation segment. The RegretNet architecture cannot handle exactly-$k$ allocations, where each buyer is allocated exactly $k$ items in the auction. The primary difference between their architecture and the RegretNet architecture lies in the allocation network. They equate the allocation problem to finding a minimum cost bipartite matching between buyers and items, for a given cost matrix. The output of the allocation network is taken as this cost matrix, the matching problem for which is solved using the Sinkhorn algorithm which takes into consideration marginal constraints such as how many items a buyer must receive. The output of this algorithm provides the probability of allocation of each item to each buyer.
@@
[RAHME22] designs a neural network architecture known as EquivarianceNet, which aims to maximize revenue while limiting DSIC violations. It overcomes three limitations of the RegretNet architecture. The first is that RegretNet cannot produce symmetric mechanisms when the optimal solution is known to be symmetric. The second limitation of RegretNet is that it requires a large number of samples to converge. EquivarianceNet makes use of domain knowledge and limits the sample space so that it can converge with fewer samples. The third limitation of RegretNet is that it cannot perform inference when the number of buyers and items is different to what it was trained on. This is overcome through the use of exchangeable matrix layers as introduced in [HARTFORD18]. The limitation in the EquivarianceNet architecture is that it assumes all buyers and items are permutation equivariant.
@@
[DUAN24] designs a revenue maximizing auction using neural networks that are inherently Dominant Strategy Incentive Compatible (DSIC) and Individually Rational (IR). Architectures like RegretNet minimize the violation of DSIC but cannot guarantee it. The authors make use of Affine Maximizer Auctions (AMAs) which are DSIC and IR. The parameters of an AMA consist of weights for each buyer and boosts associated with each allocation. The neural network architecture designed consists of an encode layer which encodes the bids into a joint representation, a menu layer to construct the allocation menu, a weight layer which learns the weights of the AMA, and a boost layer which learns the boost values of the AMA. As the mechanism is already DSIC, the only objective is to maximize the revenue.
@@
[LUONG18] proposes an optimal auction mechanism using a DL architecture for allocating edge computing resources to mobile users (i.e., miners) in a mobile blockchain environment. [QIAN19] considers Mobile Network Virtual Operator (MNVO) that provides services via an auction based NN structure to allocate resources (subchannel and power) to its users. In both these applications, the authors portray the resource allocation problem as belonging to the single item multiple buyer setting and obtain optimal auctions in the sense that they maximize the revenue of the resource provider under DSIC and IR constraints. For this, bids from users are transformed via monotone transformation functions [DUTTING23] and the second price auction with zero reserve price is used on the transformed bids to compute the allocation probability and payment rule.
@@
The authors in [ravindranath2024deepreinforcementlearningsequential] propose a revenue-optimal sequential combinatorial auction mechanism using a reinforcement learning framework. The authors introduce a fitted policy iteration approach to learn sequential auctions and implement a policy improvement step through the use of RochetNet or MenuNet. In [liu2025interpretableautomatedmechanismdesign], the authors propose an automated interpretable mechanism design using large language models and its application to both social welfare maximization & revenue maximization problems.

### v3

@@ I Introduction
How do you realize social goals in a modern society consisting of self-interested individuals or agents? This is a question facing all social planners and organisations all the time. If we solve this problem satisfactorily, it has far-reaching implications for creating robust sociological institutions, as well as, for solving numerous modern problems arising at the interface of computer science and microeconomics. Mechanism Design is an important sub-field of economic theory that strives to answer the above question. Mechanism design can be thought of as the reverse engineering of a game to achieve specific desired objectives. In mechanism design, based on the objectives to be achieved, the mechanism induces a game in a way that the desired, specific objectives are realized in an equilibrium of the game [5, 37, 36, 35]. Auctions and matching market protocols are important and popular examples of mechanisms. In this paper, we focus on auctions; however, the results presented apply to more general mechanisms as well.
@@ I-A Motivation for the Survey
Satisfying all of the properties listed above simultaneously is clearly a tall order. Mechanism design theory [36, 30, 23] is replete with celebrated results which assert that these properties cannot be simultaneously satisfied. This has led to several methodologies to surmount this limitation. One of the recently emerged, novel methodologies is an approach based on deep learning (DL). The goal of the current survey paper is to discuss this recent approach, based on Deep Learning, for designing mechanisms that achieve maximal subsets of these objectives, approximately, with as little compromise (or regret) as possible. The DL-based approach achieves this by learning a mechanism that approximately satisfies the required set of properties by minimizing a suitably defined loss function. Additionally, the DL-based approach may reduce the computational complexity of deducing the mechanism. In this paper, we present key technical details of the DL-based approach for mechanism design and provide an overview of the key methods that have emerged in this topic.
@@ I-B Outline of the Paper
The paper by Zhang [48] is the lone existing survey paper in this area. The paper by Zhang only provides an overview of how deep learning techniques are being applied to the field of online auction mechanism design. The survey explores only a few deep learning methods for designing auctions and serves only a limited purpose. Covers only papers until 2020. Our survey is comprehensive: (a) provides an overview of mechanism design with a detailed illustrative example; (b) covers four broad categories of these mechanisms: revenue maximizing (cost minimizing) mechanisms; welfare maximizing mechanisms; mechanisms with fairness in allocation; and mechanisms with budget balance; and (c) discusses three illustrative applications.
@@ II Mechanism Design: An Overview
Mechanism design can be described as the art of designing the rules of a game to achieve a specific desired outcome [36, 30]. The main focus of mechanism design is to create institutions or protocols that satisfy certain desired objectives, assuming that individual agents are rational and intelligent and will act strategically and may hold private information that is relevant to the decision at hand. In a mechanism design setting, a policy maker (also called social planner or mechanism designer) faces the problem of aggregating the announced preferences of multiple agents into a collective (or social), system-wide decision when the actual preferences of the agents are not publicly known. The essential technique that mechanism design uses to accomplish this is to induce a game among the strategic agents in such a way that in an equilibrium of the induced game, the desired system-wide solution is implemented. Informally, a mechanism induces strategic players to do what the social planner would like them to do. Mechanism design has widespread and significant applications in modern web-based and AI systems [5, 37, 36, 35].
@@ II-A Mechanisms
The notion of mechanisms was first introduced by Leonid Hurwicz (Nobel Laureate in Economic Sciences in 2007) in 1960 [36]. William Vickrey (Nobel Laureate in Economic Sciences in 1996) authored a seminal paper in 1961 [45] where he introduced the celebrated Vickrey auction (second price auction) and showed that the mechanism induces buyers in an auction to bid their true valuations without having to worry about what the other buyers bid. The Vickrey auction can be considered as a significant landmark in mechanism design theory. Edward Clarke [9] and Theodore Groves [18] came up with generalizations of Vickrey mechanisms and paved the way for the famous VCG mechanisms (Vickrey-Clarke-Groves) that define a broad class of mechanisms satisfying certain very desirable properties. The discipline of mechanism design has made phenomenal advances in the decades following the 1960s and has found widespread applicability in a wide variety of applications and areas. The scope of mechanisms includes auctions and markets, trading institutions, regulation and auditing, social choice theory, resource allocation in multi-agent systems, electronic commerce, Internet advertisement, crowdsourcing, industrial procurement, spectrum allocation, etc. [5, 37, 36, 35].
Mechanisms in which monetary transfers are allowed (for example, auctions) are known as mechanisms with money whereas mechanisms where there are no monetary transfers (such as voting mechanisms) are called mechanisms without money [36]. In this paper, we are concerned with mechanisms with money. In particular, we focus on auctions, which are arguably the most popular example of mechanisms with money. Although our discussion is centered on auctions, all the results presented here can be generalized to all mechanisms with money.
@@ II-B Auctions
Auctions provide a natural example of mechanisms with money. An auction is a mechanism for assigning a set of goods to a set of agents based on bids announced by the agents. Auctions have been used for hundreds of years for trading of goods and services and their adoption in electronic commerce has risen exponentially in the past few decades. There is intense interest in the design of auctions due to their widespread applicability to traditional settings such as paintings and memorabilia as well as their popular adoption in more recent areas such as Internet advertising, keyword auctions on search engines, airport slot allocation, spectrum allocation, industrial procurement, and resource allocation in general. A survey of auctions appears in several classic articles, including [22, 31].
@@ II-C Desirable Properties of a Mechanism
We now provide a list of important properties that a mechanism is expected to satisfy. This list is taken from [36]. As already stated, a mechanism induces a game among strategic agents in such a way that desirable properties are exhibited in an equilibrium of the game. The different equilibrium solutions that could be considered are dominant strategy equilibrium, Bayesian Nash equilibrium, and ex-post Nash equilibrium. For ease of understanding, we use the example of auctions while describing desirable properties.
@@
Individual rationality ensures that all the agents obtain non-negative utility by participating in the auction. The most powerful version of IR is ex-post IR, which implies that the utility to each participating player will be non-negative irrespective of the bids of the other players. Other forms of IR are interim IR and ex-ante IR [36, 30].
@@ II-D Possibilities and Impossibilities in Mechanism Design
The trade-offs involved in designing mechanisms with desirable properties such as incentive compatibility, individual rationality, social welfare maximization, and budget balance are quite complex and intricate [36, 30]. There is vast literature in mechanism design theory that comprises possibility results and impossibility theorems. As an example, VCG mechanisms satisfy dominant strategy incentive compatability, social welfare maximization, weak budget balance, and individual rationality. They do not satisfy strict budget balance. In fact, it is impossible for any mechanism to satisfy dominant strategy incentive compatability, social welfare maximization, and strong budget balance. Similarly, no mechanism can satisfy incentive compatability, social welfare maximization, strong budget balance, and individual rationality. When we include fairness of allocation and revenue maximization (or cost minimization), the situation gets worse and most subsets of these properties become infeasible.
@@ III An Illustrative Example: Learning a Revenue Maximizing Auction
In this section, we provide the example of an auction mechanism which is a generalization of the famous Myerson auction [34] for selling a single indivisible item. Such auctions are relevant in a wide variety of resource allocation settings. While the Myserson auction maximizes the expected revenue subject to Bayesian incentive compatibility and interim individual rationality, the extension we discuss does not satisfy these properties simultaneously and we show how the deep learning based approach is used to tackle this problem.
@@ III-B Ad-Auction: Problem Formulation
Further, designing an auction mechanism that satisfies all the desirable properties simultaneously is a challenging task and it fails in certain scenarios that are described by impossibility results [16],[2].
@@
The revelation principle [31] considers one shot revelation of true valuations of buyers to the seller by designing an IC mechanism. But it is not the case when players are selfish and may not be truthful for their personal benefits. In this scenario, a seller may be able to extract truthful valuations via multiple communications with the buyers. As opposed to the design of optimal auctions with unlimited communication [34], the authors of [33] propose a feasible auction design that maximizes the utility subject to communication feasibility constraints. Here communication feasibility specifies either delay or cost of sending the bid in terms of number of bits. As the number of items $m$ and the number of buyers $n$ increases, the cost of communicating bids or delay in communicating bids also increases. It may also take more iterative communication cycles to extract truthful information about the valuations from the buyers. The authors obtained strong results for designing optimal decentralized decision making under the absence of trade-offs between incentives and information efficiency. In [47], the authors study the relation between IC constraints and communication constraints. They also show that there exists a separation between checking IC constraints for Social Choice Function and the design of communication efficient protocol to compute the Social Choice Function for ex post Nash Implementation.
We can use the RegretNet architecture [14] to solve the $m$ items and $n$ unit demand buyers problem as shown in Fig. 1. This architecture aims to maximize the expected revenue to the seller under IC. Further details about the architecture are provided in Section IV.
Figure 1: RegretNet: The allocation network for $n$ unit demand buyers with $m$ items (Figure reproduced from [14])
@@ IV Learning Revenue Optimal Mechanisms
In [14], the authors propose an optimal auction design via RochetNet and RegretNet architectures that satisfy Dominant Strategy Incentive Compatibility (DSIC) and IR. The RochetNet architecture is based on constraining the search space so as to achieve DSIC and in this method the DSIC characterization is encoded within the Neural Network (NN) architecture. This architecture is suitable for a single buyer seeking to purchase multiple items from the auctioneer. In the case of the RegretNet architecture, approximate DSIC is achieved by differential approximation and Lagrangian augmentation. This method entails searching throughout a larger parameter space and is suitable for multi-buyer and multiple-item settings. The authors of [15] propose a DL based revenue optimal auction with private budget constraints of the buyers. This work is an extension of the RegretNet framework proposed in [14] that can handle both Bayesian IC and Conditional IC constraints. The authors in [12] introduce a new metric that quantifies Incentive Compatibility in auctions for both static and dynamic environments. This metric is data driven and is evaluated on the ad auction data.
@@
The authors in [10] design revenue maximizing auctions while approximating incentive compatibility. They build on the RegretNet architecture by introducing a differentiable matching approach in the allocation segment. The RegretNet architecture cannot handle exactly-$k$ allocations, where each buyer is allocated exactly $k$ items in the auction. The primary difference between their architecture and the RegretNet architecture lies in the allocation network. They equate the allocation problem to finding a minimum cost bipartite matching between buyers and items, for a given cost matrix. The output of the allocation network is taken as this cost matrix, the matching problem for which is solved using the Sinkhorn algorithm which takes into consideration marginal constraints such as how many items a buyer must receive. The output of this algorithm provides the probability of allocation of each item to each buyer.
@@
[40] designs a neural network architecture known as EquivarianceNet, which aims to maximize revenue while limiting DSIC violations. It overcomes three limitations of the RegretNet architecture. The first is that RegretNet cannot produce symmetric mechanisms when the optimal solution is known to be symmetric. The second limitation of RegretNet is that it requires a large number of samples to converge. EquivarianceNet makes use of domain knowledge and limits the sample space so that it can converge with fewer samples. The third limitation of RegretNet is that it cannot perform inference when the number of buyers and items is different to what it was trained on. This is overcome through the use of exchangeable matrix layers as introduced in [19]. The limitation in the EquivarianceNet architecture is that it assumes all buyers and items are permutation equivariant.
@@
[13] designs a revenue maximizing auction using neural networks that are inherently Dominant Strategy Incentive Compatible (DSIC) and Individually Rational (IR). Architectures like RegretNet minimize the violation of DSIC but cannot guarantee it. The authors make use of Affine Maximizer Auctions (AMAs) which are DSIC and IR. The parameters of an AMA consist of weights for each buyer and boosts associated with each allocation. The neural network architecture designed consists of an encode layer which encodes the bids into a joint representation, a menu layer to construct the allocation menu, a weight layer which learns the weights of the AMA, and a boost layer which learns the boost values of the AMA. As the mechanism is already DSIC, the only objective is to maximize the revenue.
@@
[28] proposes an optimal auction mechanism using a DL architecture for allocating edge computing resources to mobile users (i.e., miners) in a mobile blockchain environment. [39] considers Mobile Network Virtual Operator (MNVO) that provides services via an auction based NN structure to allocate resources (subchannel and power) to its users. In both these applications, the authors portray the resource allocation problem as belonging to the single item multiple buyer setting and obtain optimal auctions in the sense that they maximize the revenue of the resource provider under DSIC and IR constraints. For this, bids from users are transformed via monotone transformation functions [14] and the second price auction with zero reserve price is used on the transformed bids to compute the allocation probability and payment rule.
@@
The authors in [41] propose a revenue-optimal sequential combinatorial auction mechanism using a reinforcement learning framework. The authors introduce a fitted policy iteration approach to learn sequential auctions and implement a policy improvement step through the use of RochetNet or MenuNet. In [25], the authors propose an automated interpretable mechanism design using large language models and its application to both social welfare maximization & revenue maximization problems.